Circular No. 5874

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRS 1716–249 = GRO J1719–24 J. Ballet, M. Denis, M. Gilfanov, and R. Sunyaev, on behalf of the

SIGMA/GRANAT Team (Service d’Astrophysique, Centre d’Etudes de Saclay; Centre d’Etude Spatiale des Rayonnements, Toulouse; Space Research Institute, Moscow), report: “A new hard x-ray transient (GRS 1716– 249) in Ophiuchus has been discovered by SIGMA/GRANAT, during an observation of the Galactic Center region performed on Sept. 25-26. A preliminary analysis revealed the source at α = 17^{h}16^{m}40^{s}.8, δ = −24^{o}55′12′′ (equinox 1950.0; typical 90-percent confidence error circle radius of 3′). The source spectrum was found to be particularly hard, reminiscent of 1E 1740.7–2942, with flux levels of 342 ± 45, 380 ± 46, and 294 ± 85 mCrab in the energy bands 35-75, 75-150, and 150-300 keV, respectively. The source intensity remained constant during the whole observation. The source was not detected in the previous SIGMA observation of the Galactic Center region performed on Sept. 18-19, with 2σ upper limits of 50, 58, and 150 mCrab in the same three energy bands, respectively.”

B. A. Harmon, Marshall Space Flight Center (MSFC), NASA; S. N. Zhang, Universities Space Research Association; W. S. Paciesas, University of Alabama, Huntsville; and G. J. Fishman, MSFC, report for the Compton Observatory BATSE Team: “A new hard x-ray transient has been discovered using earth-occultation monitoring on Sept. 25 in the energy range 20-320 keV. Its location is α = 17^{h}19^{m}.2, δ = −24^{o}24′ (equinox 2000.0) with a 1σ error box having the following corners: α = 17^{h}24^{m}.4, δ = −26^{o}12′; 17^{h}17^{m}.6, −24^{o}30′; 17^{h}11^{m}.2, −21^{o}18′; 17^{h}20^{m}.0, −24^{o}00′. The spectrum is hard, with power law number index –2.0 in the band 20-100 keV; however, the broad-band spectrum is better fit with a thermal bremsstrahlung model with temperature kT = 70-100 keV. The temperature decreases with increasing intensity. The light curve shows the source becoming visible late on Sept. 24, with the following average intensities relative to the Crab at 20-100 keV (± 20 percent): Sept. 24, 0.07; 25, 0.46; 26, 0.82; 27, 1.1; 28, 1.2. Fast intensity variations on the order of seconds are also observed. Monitoring of the source is continuing.”

SUPERNOVA 1993J IN NGC 3031 Visual magnitude estimates (cf. IAUC 5788): Sept. 18.97 UT, 14.5 (B.

H. Granslo, Fjellhamar, Norway); 28.13, 14.3 (H. Dahle, Oslo, Norway).

1993 October 4 Daniel W. E. Green